# Quantifying the Spread of Online Incivility in Brazilian Politics

### v2

@@ Abstract
Keywords: Multidimensional Incivility |||| Brazilian Politics |||| NLP |||| Social Network |||| Social Media
1 Introduction
Social media platforms have enabled diverse political influencers—such as politicians, media outlets, journalists, and ordinary users—to produce political content and shape public opinion [19]. While social media creates an environment for ordinary users to receive political information by bypassing traditional information sources, it also provides a space for uncivil expressions and hostile attacks to unfold [23]. Previous studies have shown evidence that such uncivil behavior might reduce public political trust [43], increase polarization among groups [30], and even lead to offline violence [14]. However, incivility is not always detrimental to democracy. Studies have also shown that incivility can also draw public attention [43], encourage political participation [6], and assist marginalized groups in expressing disagreement [33]. Therefore, it is neither possible nor necessarily desirable to remove all uncivil content online [36]. This realization motivates many studies, including this one, to achieve a deeper understanding of incivility by measuring it through more nuanced categories [15, 52, 41, 5, 45].
Incivility has been previously subcategorized into behaviors such as name-calling, aspersion, vulgarity, hyperbole, and shouting/screaming notation, which involve disrespectful or rude actions towards individuals or groups [9, 44, 42]. Others argue that certain discourses may remain polite yet still be uncivil if only they violate collective norms established, such as discourse that undermines democratic values [15, 45, 49]. While many discussions about the multidimensional concept of incivility exist [41, 5], there remains a lack of relevant measurements and deeper understanding regarding who the main disseminators and audiences are and how such content spreads among users.
@@ 1 Introduction
2 Related Work
2.1 Online Incivility as a Multidimensional Concept
Incivility broadly refers to behaviors that violate collective norms and hinder cooperation among social and political agents [52]. Related concepts include impoliteness, intolerance, and toxicity. Impoliteness focuses on disrespectful communication behaviors, while intolerance—referred to by different names in various literature and defined here following [49]—highlights violations of democratic values in discourse [45, 49]. Toxicity overlaps with both concepts, emphasizing harmful elements directed at individuals or groups [2, 8, 51]. To clarify these relationships, we provide a conceptual illustration in Fig. 1
The conceptualization of (political) incivility has long been a subject of debate. For example, [16] and [9] focus on rude or disrespectful behaviors targeting individuals or groups, namely impoliteness. In contrast, [45] defines incivility as behaviors that undermine democratic values, aligning closely with intolerance as described by [49].
Recent literature has integrated these perspectives, proposing multidimensional frameworks for a more nuanced understanding of incivility [15, 41, 5]. For example, [41] distinguishes between personal-level and public-level incivility, while [5] categorizes incivility into dimensions such as impoliteness, individual delegitimization, and institutional delegitimization. Our approach summarizes and provides conceptual clarity to these various works. It stems from the collective norms that incivility violates.  [^1] The incivility dimensions in this work are part of the Swiss National Science Foundation (SNSF) project: From Uncivil Disagreement to Political Unrest? A Cross-Platform & Cross-National Analysis of the Offline Consequences of Online Incivility.
@@ 2.1 Online Incivility as a Multidimensional Concept
2.2 Quantifying Online Incivility
@@
The widespread availability of large-scale data and recent advances in machine-learning techniques have enabled the automatic detection of online incivility. Among the most commonly measured categories are toxicity and hate speech [17, 3, 50]. For instance, Google and Jigsaw developed the Perspective API, which uses machine learning to score the toxicity of input text. Similarly, hate speech, defined as ”abusive speech targeting specific group characteristics, such as ethnicity, religion, or gender”, has been tackled using various automated methods, including TF-IDF [1], lexicon-based approaches [17], deep learning methods [3], and hybrid approaches [50].
@@
Hate speech has also been distinguished from other impolite categories, such as offensive, abusive, aggressive, and cyberbullying language [12]. At a more granular level, hate speech has been subcategorized into types such as religion-based, sexist, racist, homophobic, and other forms of hate [18]. In computer science, most models for detecting hate speech are evaluated on generalized datasets [11, 37, 40]. However, researchers in other domains often fine-tune models for downstream applications for specific tasks and develop custom classifiers tailored to their unique requirements e.g., [54, 13].
@@
To predict uncivil categories, previous studies have utilized logistic regression models [54], neural networks [34], and more recently, pre-trained language models such as BERT and RoBERTa [10, 15]. For shorter social media texts, [47] introduced sentence transformers, which pool word embeddings into sentence-level representations to capture contextual meaning more effectively. However, even though classifiers for identifying toxicity and hate speech are common, classifiers targeting violations of democratic values, are rare. To the best of our knowledge, our work develops the most comprehensive classifiers, covering various dimensions of incivility.
@@
2.3 Online Political Influencers and Multi-step Flow of Uncivil Information
@@
As social media increasingly dominates communication, we have entered an era where ”any person can theoretically build an audience and grow their influence” [19]. Online influencers are particularly effective in branding and marketing due to their authentic, celebrity-like connections with broad audiences. This study focuses on political influencers, defined as influential accounts that are dedicated to promoting political stances, social identities, or preferred candidates through their content [48].
@@
Prior research identifies three key criteria for classifying political influencers: (1) their content must be political, either aggregating general political information [53] or reflecting specific political or social identities [58]; (2) their influence can be reflected by proxies such as follower counts, diffusion cascades, and potential post-earnings [4]; and (3) they leverage platform-specific features to promote content within distinct social media ecosystems [21]. Here, we identify political influencers from keywords self-disclosed in their profiles and their follower counts.
@@
Political influencers broadly include politicians, media outlets, journalists, and any influential individuals [38, 35, 32]. Among these, individual influencers play an increasingly significant role in disseminating online information [57]. These individual influencers are usually partisans, candidate supporters, and members of marginalized groups such as religious minorities, women, LGBTQ individuals, and ethnic communities.
@@
The two-step flow of communication theory highlights the role of individual opinion leaders as intermediaries, relaying information from mass media to the broader audience [28, 27]. Extensions of this model—such as horizontal two-step flows, where opinion leaders also generate their own content, and multi-step or mixed-flow frameworks, where information travels through various channels—help capture the greater complexity of modern communication. [26, 24]. Therefore, this work tests not only the two-step information flow and information from media sources, but also direct and mixed flows originating from different influential actors. Social media platforms now allow researchers to approximate these flows using platform-specific interactions. For instance, on Twitter/X, following relationships often represent direct information flows, while retweets may reflect two-step flow processes [22].
3 Data and Methods
3.1 Data Collection and Socio-political Identity Annotation
@@
We collect the Twitter accounts followed by the 271 survey participants using the Twitter API, resulting in 57,645 followers and 73,755 following pairs. We then identify political influencers based on three criteria: 1) having at least 1,000 followers, 2) profile description displaying politics-related content [29, 20], and 3) being located in Brazil (see Section 2 in Appendix for more details).
@@
3.2 Multidimension Construction and Automatic Classification of Uncivil Texts
We address RQ1 by using incivility dimensions that represent four collective norm violations, as predefined in our codebook (see Section 4 in Appendix). First, impoliteness is a typical dimension representing the violation of etiquette rules [46]. Physical harm and violent political rhetoric extend beyond the realm of etiquette and signify violations of general norm of non-violence [39]. Additionally, we consider behaviors that breach social equality & non-discrimination, and democratic norms. Hate speech and stereotyping are typical forms of uncivil language that violate social equality and non-discrimination, while violations of democratic norms involve political behaviors discrediting democratic institutions or values.
@@ 3.2 Multidimension Construction and Automatic Classification of Uncivil Texts
Step 2: Semantic Vector Projection. Annotated posts are mapped into a semantic vector space using pre-trained SentenceTransformer embeddings [47]. The centroid of the positive class is calculated as the mean of its document-level vectors. Unlabeled posts are compared to this centroid using cosine similarity to identify potential positive samples.
@@
3.3 Investigating Dynamics of Mutil-dimensional Incivility Occurrence
@@
We then identify the outliers that most significantly influence the model fitting using approximate Cook’s distance, which quantifies the influence of individual data points on the estimated regression coefficients [56]. Unlike the trends of incivility’s general ups and downs, outliers indicate specific dates when uncivil content exhibits substantial spikes.
3.4 Investigating Disseminators and Audience of Mutil-dimensional Incivility
To address RQ3, we analyze the disseminators of incivility and their audiences (survey users) across multiple dimensions. We first examine the number of survey users exposed to incivility at varying density levels. Incivility density is defined as the ratio of uncivil posts to the total number of posts by influencers. The exposed users are counted as the sum of direct followers and indirect retweets receivers of uncivil messages. To assess how incivility density influences exposure counts, we apply a quantile regression model [7]. Unlike traditional linear regression, this model estimates conditional quantiles (e.g., median or other percentiles) of the response variable, making it well-suited for our data, given the wide range of exposed audience counts at each density level. Additionally, we calculate Jaccard similarity to measure the overlap of political influencers and the survey users exposed to them across quantiles and dimensions.
@@ 3.4 Investigating Disseminators and Audience of Mutil-dimensional Incivility
3.5 Investigating Mechanisms of Information Flow of Mutil-dimensional Incivility
@@
4 Results
4.1 Performance of Automatic Classification of Uncivil Texts
@@
[…]
Etiquette Impoliteness (IMP) Messages displaying rudeness or disrespect, often using offensive language directed at individuals or groups. # Haven’t you gotten tired of the ”biggest corruption scheme in history” yet? Litany of idiots. 0.78/0.85
@@
Model performance, evaluated using ten-fold cross-validation, is summarized in Tab. 2. We calculate the metrics by aggregating all predictions from each fold to help mitigate issues caused by class imbalance within individual folds. The weighted F1 scores across dimensions range from 78% to 93%, indicating good classification performance. The trained classifiers are subsequently applied to the remaining tweets of political influencers. To ensure reliable labeling, uncivil labels - ”1” are assigned only when the predicted probabilities for the positive class are at least 0.7. The number of classified uncivil posts for each dimension is shown in Tab. 3.
@@
4.2 Different Dimensions of Incivility Break Out at Various Stages
@@
4.3 Dominant Disseminators Differ but Share Similar Identities Across Dimensions
@@
[…]
black Dimension (DIM) G-test Chi2 (AT) $p$-value (AT) $D_{\text{IMP}}$ $176.88$ $1.56\cdot 10^{-35}$ $D_{\text{PHAVPR}}$ $59.54$ $5.57\cdot 10^{-11}$ $D_{\text{HSST}}$ $130.58$ $9.69\cdot 10^{-26}$ $D_{\text{THREAT}}$ $12.39$ $5.38\cdot 10^{-2}$ Dimension (DIM) G-test Chi2 (IE) $p$-value (IE) $D_{\text{IMP}}$ $115.02$ $2.38\cdot 10^{-2}$ $D_{\text{PHAVPR}}$ $134.73$ $3.72\cdot 10^{-4}$ $D_{\text{HSST}}$ $95.64$ $1.81\cdot 10^{-1}$ $D_{\text{THREAT}}$ $148.10$ $7.86\cdot 10^{-6}$
@@
4.4 Direct Information Flow Dominates in Uncivil Communication Online, Whereas Mixed Information Flow is Significant
@@
[…]
black # $D_{\text{IMP}}$ L $D_{\text{PHAVPR}}$ L $D_{\text{HSST}}$ L $D_{\text{THREAT}}$ L 1 An** PO Ha** PO Ga** PO Gl** MO 2 Lu** PO An** PO An** PO UO** MO 3 La** IN An** JL Ja** PO Ca** JL 4 Gl** PO Mo** JL Tr** JL Gl** PO 5 Ca** PO Gi** JL NP** IN Re** JL 6 Ha** PO Fo** MO Bo** PO Gu** JL 7 De** IN Xi** JL La** JL Me** MO 8 Jo** PO UO** MO Re** JL EC** JL 9 Ja** PO Er** IN Se** PO Es** MO 10 Fl** PO Th** IN Fo** MO Is** IN

### v3

@@ Abstract
Introduction
Social media platforms have enabled diverse political influencers—such as politicians, media outlets, journalists, and ordinary users—to produce political content and shape public opinion (Goodwin et al. 2023). While social media creates an environment for ordinary users to receive political information by bypassing traditional information sources, it also provides a space for uncivil expressions and hostile attacks to unfold (Heseltine and Dorsey 2022). Previous studies have shown evidence that such uncivil behavior might reduce public political trust (Mutz and Reeves 2005), increase polarization among groups (Kim and Kim 2019), and even lead to offline violence (Gallacher, Heerdink, and Hewstone 2021). However, incivility is not always detrimental to democracy. Studies have also shown that incivility can also draw public attention (Mutz and Reeves 2005), encourage political participation (Brooks and Geer 2007), and assist marginalized groups in expressing disagreement (Lozano-Reich and Cloud 2009). Therefore, it is neither possible nor necessarily desirable to remove all uncivil content online (Masullo Chen et al. 2019). This realization motivates many studies, including this one, to achieve a deeper understanding of incivility by measuring it through more nuanced categories (Gao et al. 2024; Stryker, Conway, and Danielson 2016; Muddiman 2017; Bentivegna and Rega 2022; Papacharissi 2004).
Incivility has been previously subcategorized into behaviors such as name-calling, aspersion, vulgarity, hyperbole, and shouting/screaming notation, which involve disrespectful or rude actions towards individuals or groups (Coe, Kenski, and Rains 2014; Otto, Lecheler, and Schuck 2020; Mutz 2007). Others argue that certain discourses may remain polite yet still be uncivil if only they violate collective norms established, such as discourse that undermines democratic values (Gao et al. 2024; Papacharissi 2004; Rossini 2022). While many discussions about the multidimensional concept of incivility exist (Muddiman 2017; Bentivegna and Rega 2022), there remains a lack of relevant measurements and deeper understanding regarding who the main disseminators and audiences are and how such content spreads among users.
@@ Introduction
Related Work
Online Incivility as a Multidimensional Concept
Incivility broadly refers to behaviors that violate collective norms and hinder cooperation among social and political agents (Stryker, Conway, and Danielson 2016). Related concepts include impoliteness, intolerance, and toxicity. Impoliteness focuses on disrespectful communication behaviors, while intolerance—referred to by different names in various literature and defined here following Rossini—highlights violations of democratic values in discourse (Papacharissi 2004; Rossini 2022). Toxicity overlaps with both concepts, emphasizing harmful elements directed at individuals or groups (Aleksandric et al. 2024; Chong and Kwak 2022; Singh et al. 2024). To clarify these relationships, we provide a conceptual illustration in Fig. 1
The conceptualization of (political) incivility has long been a subject of debate. For example, (Gervais 2015) and (Coe, Kenski, and Rains 2014) focus on rude or disrespectful behaviors targeting individuals or groups, namely impoliteness. In contrast, (Papacharissi 2004) defines incivility as behaviors that undermine democratic values, aligning closely with intolerance as described by (Rossini 2022).
Recent literature has integrated these perspectives, proposing multidimensional frameworks for a more nuanced understanding of incivility (Gao et al. 2024; Muddiman 2017; Bentivegna and Rega 2022). For example, (Muddiman 2017) distinguishes between personal-level and public-level incivility, while (Bentivegna and Rega 2022) categorizes incivility into dimensions such as impoliteness, individual delegitimization, and institutional delegitimization. Our approach summarizes and provides conceptual clarity to these various works. It stems from the collective norms that incivility violates.  [^1] The incivility dimensions in this work are part of the Swiss National Science Foundation (SNSF) project: From Uncivil Disagreement to Political Unrest? A Cross-Platform & Cross-National Analysis of the Offline Consequences of Online Incivility.
@@
Quantifying Online Incivility
@@
The widespread availability of large-scale data and recent advances in machine-learning techniques have enabled the automatic detection of online incivility. Among the most commonly measured categories are toxicity and hate speech (Gitari et al. 2015; Badjatiya et al. 2017; Shah et al. 2021). For instance, Google and Jigsaw developed the Perspective API, which uses machine learning to score the toxicity of input text. Similarly, hate speech, defined as "abusive speech targeting specific group characteristics, such as ethnicity, religion, or gender", has been tackled using various automated methods, including TF-IDF (Akuma, Lubem, and Adom 2022), lexicon-based approaches (Gitari et al. 2015), deep learning methods (Badjatiya et al. 2017), and hybrid approaches (Shah et al. 2021).
@@
Hate speech has also been distinguished from other impolite categories, such as offensive, abusive, aggressive, and cyberbullying language (Founta et al. 2018). At a more granular level, hate speech has been subcategorized into types such as religion-based, sexist, racist, homophobic, and other forms of hate (Gomez et al. 2020). In computer science, most models for detecting hate speech are evaluated on generalized datasets (Davidson et al. 2017; Mathew et al. 2021; Mollas et al. 2022). However, researchers in other domains often fine-tune models for downstream applications for specific tasks and develop custom classifiers tailored to their unique requirements e.g., (Theocharis et al. 2020; Frimer et al. 2023).
@@
To predict uncivil categories, previous studies have utilized logistic regression models (Theocharis et al. 2020), neural networks (Maity et al. 2018), and more recently, pre-trained language models such as BERT and RoBERTa (Davidson, Sun, and Wojcieszak 2020; Gao et al. 2024). For shorter social media texts, (Reimers and Gurevych 2019) introduced sentence transformers, which pool word embeddings into sentence-level representations to capture contextual meaning more effectively. However, even though classifiers for identifying toxicity and hate speech are common, classifiers targeting violations of democratic values, are rare. To the best of our knowledge, our work develops the most comprehensive classifiers, covering various dimensions of incivility.
@@
Online Political Influencers and Multi-step Flow of Uncivil Information
@@
As social media increasingly dominates communication, we have entered an era where "any person can theoretically build an audience and grow their influence" (Goodwin et al. 2023). Online influencers are particularly effective in branding and marketing due to their authentic, celebrity-like connections with broad audiences. This study focuses on political influencers, defined as influential accounts that are dedicated to promoting political stances, social identities, or preferred candidates through their content (Riedl, Lukito, and Woolley 2023).
@@
Prior research identifies three key criteria for classifying political influencers: (1) their content must be political, either aggregating general political information (Su and Marbach 2024) or reflecting specific political or social identities (Zhang, Chen, and Lukito 2023); (2) their influence can be reflected by proxies such as follower counts, diffusion cascades, and potential post-earnings (Bakshy et al. 2011); and (3) they leverage platform-specific features to promote content within distinct social media ecosystems (Harris, Foxman, and Partin 2023). Here, we identify political influencers from keywords self-disclosed in their profiles and their follower counts.
@@
Political influencers broadly include politicians, media outlets, journalists, and any influential individuals (McGregor 2020; Marozzo and Bessi 2018; Lee 2015). Among these, individual influencers play an increasingly significant role in disseminating online information (Wu et al. 2011). These individual influencers are usually partisans, candidate supporters, and members of marginalized groups such as religious minorities, women, LGBTQ individuals, and ethnic communities.
@@
The two-step flow of communication theory highlights the role of individual opinion leaders as intermediaries, relaying information from mass media to the broader audience (Katz and Lazarsfeld 1955; Katz 1957). Extensions of this model—such as horizontal two-step flows, where opinion leaders also generate their own content, and multi-step or mixed-flow frameworks, where information travels through various channels—help capture the greater complexity of modern communication. (Hunt and Gruszczynski 2024; Hilbert et al. 2017). Therefore, this work tests not only the two-step information flow and information from media sources, but also direct and mixed flows originating from different influential actors. Social media platforms now allow researchers to approximate these flows using platform-specific interactions. For instance, on Twitter/X, following relationships often represent direct information flows, while retweets may reflect two-step flow processes (Hemsley 2019).
Data and Methods
Data Collection and Socio-political Identity Annotation
@@
We collect the Twitter accounts followed by the 271 survey participants using the Twitter API, resulting in 57,645 followers and 73,755 following pairs. We then identify political influencers based on three criteria: 1) having at least 1,000 followers, 2) profile description displaying politics-related content (Khamis, Ang, and Welling 2017; Harff and Schmuck 2023), and 3) being located in Brazil (see Section 2 in Appendix for more details).
@@
Multidimension Construction and Automatic Classification of Uncivil Texts
We address RQ1 by using incivility dimensions that represent four collective norm violations, as predefined in our codebook (see Section 4 in Appendix). First, impoliteness is a typical dimension representing the violation of etiquette rules (Rega, Marchetti, and Stanziano 2023). Physical harm and violent political rhetoric extend beyond the realm of etiquette and signify violations of general norm of non-violence (Miller 1984). Additionally, we consider behaviors that breach social equality & non-discrimination, and democratic norms. Hate speech and stereotyping are typical forms of uncivil language that violate social equality and non-discrimination, while violations of democratic norms involve political behaviors discrediting democratic institutions or values.
@@
Step 2: Semantic Vector Projection. Annotated posts are mapped into a semantic vector space using pre-trained SentenceTransformer embeddings (Reimers and Gurevych 2019). The centroid of the positive class is calculated as the mean of its document-level vectors. Unlabeled posts are compared to this centroid using cosine similarity to identify potential positive samples.
@@
Investigating Dynamics of Mutil-dimensional Incivility Occurrence
@@
We then identify the outliers that most significantly influence the model fitting using approximate Cook’s distance, which quantifies the influence of individual data points on the estimated regression coefficients (Wood 2017). Unlike the trends of incivility’s general ups and downs, outliers indicate specific dates when uncivil content exhibits substantial spikes.
Investigating Disseminators and Audience of Mutil-dimensional Incivility
To address RQ3, we analyze the disseminators of incivility and their audiences (survey users) across multiple dimensions. We first examine the number of survey users exposed to incivility at varying density levels. Incivility density is defined as the ratio of uncivil posts to the total number of posts by influencers. The exposed users are counted as the sum of direct followers and indirect retweets receivers of uncivil messages. To assess how incivility density influences exposure counts, we apply a quantile regression model (Buchinsky 1998). Unlike traditional linear regression, this model estimates conditional quantiles (e.g., median or other percentiles) of the response variable, making it well-suited for our data, given the wide range of exposed audience counts at each density level. Additionally, we calculate Jaccard similarity to measure the overlap of political influencers and the survey users exposed to them across quantiles and dimensions.
@@
Investigating Mechanisms of Information Flow of Mutil-dimensional Incivility
@@
Results
Performance of Automatic Classification of Uncivil Texts
@@ Results
Related Norm Dimension Definition Example (translated) Gwet’s AC2 Score/Agreement
Etiquette Impoliteness (IMP) Messages displaying rudeness or disrespect, often using offensive language directed at individuals or groups. # Haven’t you gotten tired of the "biggest corruption scheme in history" yet? Litany of idiots. 0.78/0.85
@@
Model performance, evaluated using ten-fold cross-validation, is summarized in Tab. 2. We calculate the metrics by aggregating all predictions from each fold to help mitigate issues caused by class imbalance within individual folds. The weighted F1 scores across dimensions range from 78% to 93%, indicating good classification performance. The trained classifiers are subsequently applied to the remaining tweets of political influencers. To ensure reliable labeling, uncivil labels - "1" are assigned only when the predicted probabilities for the positive class are at least 0.7. The number of classified uncivil posts for each dimension is shown in Tab. 3.
@@
Different Dimensions of Incivility Break Out at Various Stages
@@
Dominant Disseminators Differ but Share Similar Identities Across Dimensions
@@
Dimension (DIM) G-test Chi2 (AT) $p$-value (AT)
$D_{\text{IMP}}$ $176.88$ $1.56\cdot 10^{-35}$
$D_{\text{PHAVPR}}$ $59.54$ $5.57\cdot 10^{-11}$
$D_{\text{HSST}}$ $130.58$ $9.69\cdot 10^{-26}$
$D_{\text{THREAT}}$ $12.39$ $5.38\cdot 10^{-2}$
Dimension (DIM) G-test Chi2 (IE) $p$-value (IE)
$D_{\text{IMP}}$ $115.02$ $2.38\cdot 10^{-2}$
$D_{\text{PHAVPR}}$ $134.73$ $3.72\cdot 10^{-4}$
$D_{\text{HSST}}$ $95.64$ $1.81\cdot 10^{-1}$
$D_{\text{THREAT}}$ $148.10$ $7.86\cdot 10^{-6}$
@@
Direct Information Flow Dominates in Uncivil Communication Online, Whereas Mixed Information Flow is Significant
@@
# $D_{\text{IMP}}$ L $D_{\text{PHAVPR}}$ L $D_{\text{HSST}}$ L $D_{\text{THREAT}}$ L
1 An** PO Ha** PO Ga** PO Gl** MO
2 Lu** PO An** PO An** PO UO** MO
3 La** IN An** JL Ja** PO Ca** JL
4 Gl** PO Mo** JL Tr** JL Gl** PO
5 Ca** PO Gi** JL NP** IN Re** JL
6 Ha** PO Fo** MO Bo** PO Gu** JL
7 De** IN Xi** JL La** JL Me** MO
8 Jo** PO UO** MO Re** JL EC** JL
9 Ja** PO Er** IN Se** PO Es** MO
10 Fl** PO Th** IN Fo** MO Is** IN